### v1

@@ Abstract
Dense features are important for detecting minute objects in images. Unfortunately, despite the remarkable efficacy of the CNN models in multi-scale object detection, CNN models often fail to detect smaller objects in images due to the loss of dense features during the pooling process. Atrous convolution addresses this issue by applying sparse kernels. However, sparse kernels often can lose the multi-scale detection efficacy of the CNN model. In this paper, we propose an object detection model using a Switchable (adaptive) Atrous Convolutional Network (SAC-Net) based on the efficientDet model. A fixed atrous rate limits the performance of the CNN models in the convolutional layers. To overcome this limitation, we introduce a switchable mechanism that allows for dynamically adjusting the atrous rate during the forward pass. The proposed SAC-Net encapsulates the benefits of both low-level and high-level features to achieve improved performance on multi-scale object detection tasks, without losing the dense features. Further, we apply a depth-wise switchable atrous rate to the proposed network, to improve the scale-invariant features. Finally, we apply global context on the proposed model. Our extensive experiments on benchmark datasets demonstrate that the proposed SAC-Net outperforms the state-of-the-art models by a significant margin in terms of accuracy. The codes are available at
@@ 1 Introduction
In this study, we applied the concept of switchable atrous convolution (motivated by detectors ), on a lighter model called EfficientDet efficientdet , to reduce the GPU overhead and FLOPs for the proposed method. Further, we apply the switchable atrous rates depthwise, at the backbone model, to enhance the effect of atrous convolution in the model accuracy. Finally, we apply global context before and after the depthwise convolution layers, to make the model scale invariant. The contributions of this study can be summarized as follows:
  • We apply an adaptive atrous Conv layer with different atrous rates, capable of adapting with varying depths of the architecture.
@@ 3.4 Depthwise switchable atrous Conv layer with the different atrous rate (DSAC)
Figure 2: Depthwise switchable atrous Conv layer with different atrous(DSAC).Switchable Atrous Convolution (DSAC). We convert each $3\times 3$ convolutional layer in the baseline Efficientnet to DSAC, which gradually alternates the atrous rates used for convolutional computation. Two global context modules add image-level information to the features.
@@ 3.5 Depthwise atrous with pointwise switchable Conv layer (DAPSC)
Combining Equations (4) and (5) with a switch function results in the following:
@@
Figure 3: Depthwise atrous with pointwise switchable Conv layer(DAPSC).Switchable Atrous Convolution (DSAPSC). We convert each 3x3 convolutional layer in the baseline Efficientnet to DAC, which gradually alternates the atrous rates used for convolutional computation and added a pointwise switch function(PSC). Two global context modules add image-level information to the features.
In Equation (4), the input feature map $x$ is first passed through a DConv layer with an atrous rate 1, denoted as $DConv(x,w,1)$. The output of the DConv layer is then passed through a squeeze-and-excitation (SE) layer, denoted as $SE(x)$, which adaptively recalibrates the feature responses based on channel-wise information. Finally, the output of the SE layer is passed through a point-wise convolutional layer, denoted as $PConv(x)$, with an atrous rate of 1.
Similarly, in Equation (5), the input feature map $x$ is passed through a DConv layer with atrous rate 3, denoted as $DConv(x,w,3)$. The output of the DConv layer is then passed through a SE layer, denoted as $SE(x)$, and finally, through a PConv layer with an atrous rate of 1.
The combination of Equations (4) and (5) with a switch function ($S(x)$) results in the following operation (6): for each location in the feature map, the switch function generates a binary mask that decides which operation to perform - either pass the feature map through $PConv(x,w,1)$ or $PConv(x,w,r)$ (with $r=3$). The binary mask is then used to select either the output of $PConv(x,w,1)$ or the output of $PConv(x,w,r)$ for each location in the feature map, and the selected feature maps are combined using element-wise addition to produce the final output feature map.
@@ 3.6 Global context before and after the depthwise convolution layers
Two global context modules are inserted before and after the depthwise separable Conv layer’s primary component as shown in Fig.4. These two modules are lightweight because a global average pooling layer first compresses the input characteristics. The results are incorporated back into the mainstream. We observe that the detection performance is improved by including the global context data before the depthwise separable Conv layers $DConv(x,w,1)$, Where $x$ is input, $w$ is the weight, and 1 is atrous rate. The performances of depth-wise separable Conv layers with and without the global context modules are shown in the ablation research.
$DConv(x,w,1)\longrightarrow PrG(x)+DConv(x,w,1)+PoG(x),$ (7)
@@
The combined feature map produced by Equation (7) contains local and global context information.
@@ 5.1 Ablation Studies
D0(Original) 1 .00001 SGD 16 31.5
D1(Original) 1 .00001 SGD 8 36.3
D2(Original) 1 .00001 SGD 4 39.0
D0+Global context 2 .0001 SGD 16 32.2
D1+Global context 2 .0001 SGD 8 37.8
D2+Global context 2 .0001 SGD 4 40.7
D0+DSAC 3 .00001 SGd 8 31.6
D1+DSAC 3 .00001 SGD 8 38. 0
D2+DSAC 3 .00001 SGD 4 40.6
D0+DSAC+Global Context 2 .00001 SGD 16 32.5
D1+DSAC+Global Context 2 .00001 SGD 4 38.8
D2+DSAC+Global Context 2 .00001 SGD 4 40.7
D0+DAPSC 2 .00001 SGD 16 32.3
D1+DAPSC 2 .00001 SGD 4 38.1
D2+DAPSC 2 .00001 SGD 4 41.0
D0+DAPSC+ Global Context 2 .00001 SGD 16 32.4
D1+DAPSC+ Global Context 2 .00001 SGD 4 37.5
D2+DAPSC+ Global Context 2 .00001 SGD 4 40.6
@@ 6 Conclusion
A depthwise switchable atrous convolutional network is proposed in this study. The proposed model uses a switchable mechanism to control the use of different rates of atrous (dilated) convolution operations, including traditional atrous convolution, depthwise atrous convolution, and pointwise atrous convolution. The idea behind the depthwise switchable atrous convolutional network is to allow the network to automatically switch between these different types of atrous convolution operations based on the input data so that it can be useful for detecting objects that appear in different scales in the images, and improve the accuracy of the proposed method. The proposed scale-invariant feature can be extended to work on videos, to track objects across frames with varying scales. Further, the proposed atrous convolution can be tested with the YOLO family of object detectors, to analyze the effect.

### v2

@@ Abstract
Dense features are important for detecting differnet scale objects in images. Unfortunately, despite of the remarkable efficacy of the CNN models in multi-scale object detection, CNN models often fail to detect different scale objects in images due to using similar type of CNN features . Atrous convolution address this issue by applying sparse kernels. However, sparse kernels often can lose the multi-scale detection efficacy of the CNN model. In this paper, we propose an object detection model using a Switchable Atrous Convolutional Network (SAC-Net) based on the efficientDet model. A fixed atrous rate limits the performance of the CNN models in the convolutional layers. To overcome this limitation, we introduce a switchable mechanism that allows for dynamically adjusting the atrous rate during the forward pass. The proposed SAC-Net encapsulates the benefits of both low-level and high-level features to achieve improved performance on multi-scale object detection tasks, without losing the dense features. Further, we apply a depth-wise switchable atrous rate to the proposed network, to improve the scale-invariant features. Finally, we apply global context on the proposed model. Our extensive experiments on benchmark datasets demonstrate that the proposed SAC-Net outperforms the state-of-the-art models by a significant margin in terms of accuracy. The codes are available at
@@ 1 Introduction
In this study, we applied the concept of switchable atrous convolution (motivated by detectors ), on a lighter model called EfficientDet efficientdet , to reduce the GPU overhead and FLOPs for the proposed method. Further, we apply the switchable atrous rates depthwise, at the backbone model, to enhance the effect of atrous convolution in the model accuracy. Finally, we apply global context before and after the depthwise convolution layers, to make the model scale invariant. The contributions of this study is summarized as follows:
  • In this proposed work using an adaptive atrous Conv layer with different atrous rates, capable of adapting with varying depths of the architecture.
@@ 3.4 Depthwise switchable atrous Conv layer with the different atrous rate (DSAC)
Figure 2: Depthwise switchable atrous Conv layer with different atrous(DSAC). We convert each $3\times 3$ convolutional layer in the baseline Efficientnet to DSAC, which gradually alternates the atrous rates used for convolutional computation. Two global context modules add image-level information to the features.
@@ 3.5 Depthwise atrous with pointwise switchable Conv layer (DAPSC)
Combining (4) and (5) with a switch function $S(x)$ results in the following:
@@
Figure 3: Depthwise atrous with pointwise switchable Conv layer(DAPSC) . We convert each $3\times 3$ convolutional layer in the baseline Efficientnet to DAC, which gradually alternates the atrous rates used for convolutional computation and added a pointwise switch function(PSC). Two global context modules add image-level information to the features.
In (4), the input feature map $x$ is first passed through a DConv layer with an atrous rate 1, denoted as $DConv(x,w,1)$. The output of the DConv layer is then passed through a squeeze-and-excitation (SE) layer, denoted as $SE(x)$, which adaptively recalibrates the feature responses based on channel-wise information. Finally, the output of the SE layer is passed through a point-wise convolutional layer, denoted as $PConv(x)$, with an atrous rate of 1.
Similarly, in (5), the input feature map $x$ is passed through a DConv layer with atrous rate 3, denoted as $DConv(x,w,3)$. The output of the DConv layer is then passed through a SE layer, denoted as $SE(x)$, and finally, through a PConv layer with an atrous rate of 1.
The combination of (4) and (5) with a switch function ($S(x)$) results in the following operation (6): for each location in the feature map, the switch function generates a binary mask that decides which operation to perform - either pass the feature map through $PConv(x,w,1)$ or $PConv(x,w,r)$ (with $r=3$). The binary mask is then used to select either the output of $PConv(x,w,1)$ or the output of $PConv(x,w,r)$ for each location in the feature map, and the selected feature maps are combined using element-wise addition to produce the final output feature map.
@@ 3.6 Global context before and after the depthwise convolution layers
Two global context modules are inserted before and after the depthwise separable Conv layer’s primary component as shown in Fig.4. These two modules are lightweight because a global average pooling layer first compresses the input characteristics. The results are incorporated back into the mainstream. We observe that the detection performance is improved by including the global context data before the depthwise separable Conv layers $DConv(x,w,1)$, Where, $x$ , $w$ and 1, are input weight and atrous rate. The performances of depth-wise separable Conv layers with and without the global context modules are shown in the ablation research.
$DConv(x,w,1)\longrightarrow PrG(x)+DConv(x,w,1)+PoG(x)$ (7)
@@
The combined feature map produced by (7) contains local and global context information.
@@ 5.1 Ablation Studies
D0(Original) 1 0.00001 SGD 16 31.5
D1(Original) 1 0.00001 SGD 8 36.3
D2(Original) 1 0.00001 SGD 4 39.0
D0+Global context 2 0.0001 SGD 16 32.2
D1+Global context 2 0.0001 SGD 8 37.8
D2+Global context 2 0.0001 SGD 4 40.7
D0+DSAC 3 0.00001 SGd 8 31.6
D1+DSAC 3 0.00001 SGD 8 38. 0
D2+DSAC 3 0.00001 SGD 4 40.6
D0+DSAC+Global Context 2 0.00001 SGD 16 32.5
D1+DSAC+Global Context 2 0.00001 SGD 4 38.8
D2+DSAC+Global Context 2 0.00001 SGD 4 40.7
D0+DAPSC 2 0.00001 SGD 16 32.3
D1+DAPSC 2 0.00001 SGD 4 38.1
D2+DAPSC 2 0.00001 SGD 4 41.0
D0+DAPSC+ Global Context 2 0.00001 SGD 16 32.4
D1+DAPSC+ Global Context 2 0.00001 SGD 4 37.5
D2+DAPSC+ Global Context 2 0.00001 SGD 4 40.6
@@ 6 Conclusion
A depthwise switchable atrous convolutional network is proposed in this research work. The proposed model equipped a switchable mechanism to control the use of different rates of atrous (dilated) convolution operations, including traditional atrous convolution, depthwise atrous convolution, and pointwise atrous convolution. The idea behind the depthwise switchable atrous convolutional network is to allow the network to automatically switch between these different types of atrous convolution operations based on the input data so that it can be useful for detecting objects that appear in different scales in the images, and improve the accuracy of the proposed method. The proposed scale-invariant feature can be extended to work on videos, to track objects across frames with varying scales. Further, the proposed atrous convolution can be tested with the YOLO family of object detectors, to analyze the effect.